Joe inherited some money from his father. He invested one-half of this amount at 5%, one-third (1/3) of this amount at 6%, and the rest at 3%. If the total income from these investments was $300,
What was the amount he inherited???

Answer by mananth(16949)

Joe inherited some money from his father. He invested one-half of this amount at 5%, one-third (1/3) of this amount at 6%, and the rest at 3%. If the total income from these investments was $300,
What was the amount he inherited???
...
let Joe inherit $x
x/2 @ 5% interest =
x/3 @ 6% interest = .
x/2+x/3 = 5x/6 is total on two investments
balance = x-5x/6
=x/6
(x/2) * 5% +(x/3)*6% +(x/6)*3%=300
0.025x+0.02x+0.005x=300
o.05x=300
x=300/0.05
x= $6000 is the amount inherited
...
m.ananth@hotmail.ca